How much work is needed to move a proton from a point with an potential of +75 V to a point where it is -25 V?

1
Expert's answer
2022-05-18T08:31:17-0400
"W=q(V_2-V_1)\\
=1.6*10^{-19}*(75-(-25))=1.6*10^{-17}\rm \: J"